A box contains 10 white and 20 black marbles one selects at random 10 with replacement compute the probability of obtaining at least 2 white marbles
1
Expert's answer
2020-04-06T18:01:24-0400

Probability of picking a white ball with replacement is "\frac{10}{30}=\frac{1}{3}" . Picking balls ten time with replacement forms a Binomial distribution.

"P(\ge2)=1-P(\le1)"

"P(=0)=(\frac{2}{3})^{10}=0.0173415"

"P(=1)=\frac{10}{3}×(\frac{2}{3})^9=0.08671"

"P(\u22652)=1-0.0173415-0.008671=0.895949"

Similarly the Excel formula, =1-BINOM.DIST(1,10,1/3,TRUE) can be used to obtain the probability.